Light with wavelength 405nm incidents on a diffraction grating with 1000 lines per mm. At what angle is the 1st order diffracted beam found?
Answer #1

Solution given Np = 1000 lines Imm 10-3m = 10-6m 1000 d = 1 Hg 405X10-9m d-4050m the condition for maxima dsino = ml first or